# Unifying the Extremes: Developing a Unified Model for Detecting and Predicting Extremist Traits and Radicalization

### v1

@@ 7.3 User-level Analysis
Figure 6: Incel User Case Study: A timeline illustrating the phases of incel extremist community member exhibiting changes in extremist beliefs. Five months prior to joining the extremist community, the user exhibited mild anxiety and frustration with dating. Shortly after joining, they show signs of resentment over rejection and envy, while implying a perceived benefit of joining (to help rid some misery). At 3-4 months, they begin to blame their circumstances on society and use increasingly dehumanizing language about women; note the progression from girl, females, to femoids. At 6-10 months, the user’s rhetoric becomes retributive and violent. The spans are paraphrased extracts from the user’s posts.
Using the same dataset, we additionally examined whether engaging in such communities amplifies extremist traits. We perform local polynomial regression (LOESS) on the extremism scores for each of the users joining the incel community over a period and the users joining a random community over a period of 24 months to derive smooth trajectories of extremism scores. Figure 5 compares the extremism scores of users in incel forums to those in general Reddit forums, analyzing changes in scores before and after joining a new group. For users joining a random, non-extremist forum, we found no significant change in their extremist trait scores. However, for those joining the incel community, we observed a noticeable increase in these traits leading up to and following their entry into the forum; Fig.  6 analyzes the progression of one user’s extremism over this timeline. Despite this, the increase was not as dramatic as one might expect, which aligns with recent research questioning the validity of the “echo chamber” metaphor in explaining polarization and radicalization (e.g., Törnberg and Törnberg 2024). This suggests that joining an extremist group may not automatically intensify extremist tendencies, challenging the assumption that such communities solely function as amplifiers of radicalization.
@@ 9 Limitations and Future Work
Appendix A Appendix
A.1 Ethics Statement
This study involves the analysis of online data from extremist and general social media forums, raising ethical considerations related to privacy, consent, and the potential impact of the research. All data used in this study were collected from publicly available sources where users posted content in public forums. No private or personally identifiable information was accessed or used in the analysis. In order to mitigate any potential harm, we took careful steps to anonymize the data, and user-level analyses were conducted with caution, ensuring that no individual participants could be identified.
The potential risks of using this data for research include the reinforcement of stigmas surrounding certain groups, as well as unintended consequences that may arise from the misuse of findings. To address this, we approached the analysis from a neutral, empirical perspective, aiming to better understand patterns of radicalization while being mindful of the broader social context in which this information is used.
Additionally, we acknowledge that studying extremist discourse may inadvertently provide insights into how these groups operate, which could be used for harmful purposes. To prevent this, we focused on developing models that contribute to early detection and intervention strategies, with the aim of mitigating the harm caused by radicalization. We are committed to ensuring that the findings of this work are used to support efforts in counter-radicalization, public safety, and mental health intervention.
This research was conducted in line with institutional guidelines for the ethical use of publicly available data and followed the principles of responsible AI development, particularly regarding the societal impact of extremism detection and intervention strategies.
A.2 Compute and Hyperparameters
@@
A.3 Term Definitions and Disambiguation
Extremism is the promotion or advancement of an ideology based on violence, hatred, or intolerance that aims to negate fundamental rights, undermine democratic systems, or create permissive environments for such actions  (Government of United Kingdom 2024). It represents views and behaviors that are not just unconventional, but rather far removed from mainstream societal norms and attitudes, advocating for extreme measures that threaten democratic systems and ability of people to live equally under the law. The manifestation of extremist beliefs spans a spectrum, ranging from subtle expressions of intolerance to overt displays of hatred, and ultimately culminating in explicit acts of violence against marginalized groups  (Schmid 2022; Knight, Woodward, and Lancaster 2017).
Radicalization is the process of change by which an individual or group comes to adopt increasingly extreme views in opposition to a political, social, economic or religious status quo  (Police 2009) . This could result in explicit outcomes such as violence or hate-mongering, leading to degeneration of the individual or the group into extremist ideologies. Radicalization and extremist tendencies usually go hand-in-hand, however, while radicalization focuses more on developing beliefs advocating for drastic transformation of the social structure, tending towards more and more extreme measures, while extremism comprises of not just the ideologies, but intentional actions and behaviors to propagate those ideologies as well  (Trip et al. 2019).
Hate speech is public speech or texts that encourage violence against specific social groups based on race, ethnicity, religion etc.  (Dictionary 2024). While hate speech is commonly studied in NLP contexts  (MacAvaney et al. 2019), studying extremism and radicalization online goes beyond just looking for hate speech or speech inciting violence. Although hate speech may stem from deep-rooted extremist beliefs (Bilewicz and Soral 2020), it can also manifest from targeted hostility toward individuals based on personal grievances unrelated to their social identity.  (Fortuna, Soler, and Wanner 2020; Piot, Martín-Rodilla, and Parapar 2024)This distinction highlights that hate speech, while potentially overlapping with extremist ideologies and radicalization processes, does not necessarily indicate their presence. Our method is comprehensive – it goes beyond language-based violence incitement to examine fundamental aspects such as underlying beliefs, documented belief transitions, personality traits, dogmatic tendencies, and political, religious, and economic orientations, all grounded in extensive academic research to accurately identify extremist inclinations and radicalization patterns.
A.4 Banned and Not-Banned Subreddit Samples
Banned r/Right_Wing_Politics r/europeannationalism r/Identitarians r/DebateAltRight r/AltRightChristian r/pol r/NationalSocialism r/AgainstTheChimpire r/donaldtrump Not-Banned r/AmericanPolitics r/Classical_Liberals r/gunpolitics r/Ultraleft r/CanadianPolitics r/PeoplesPartyofCanada r/JordanPeterson r/justicedemocrats r/StillSandersForPres r/CapitalismVSocialism
A.5 Additional Figures
Figure A1: Early detection power of the Extremist Eleven scores for identifying whether users will actively engage with the incel community in the future. The shaded region represents the 95% confidence interval of the predicted AUC scores. We find that the prediction is already better than chance 8-10 months before a user joins the incel forum.
@@ A.5 Additional Figures
The ?old-fashioned ways? and ?old-fashioned values? still show the best way to live. 0.72 Right-Wing Authoritarianism
@@
God?s laws about abortion, pornography and marriage must be strictly followed before it is too late, violations must be punished. 0.79 Right-Wing Authoritarianism
@@
God?s laws about abortion, pornography and marriage must be strictly followed before it is too late, violations must be punished. 0.79 Right-Wing Authoritarianism
@@
Table A1: Factors of the “Extremist Eleven” along with each factor’s highest scoring extremism scale items.

### v2

@@ 7 Results and Discussion
Figure 6: Early detection power of the Extremist Eleven scores for identifying whether users will actively engage with the incel community in the future. The shaded region represents the 95% confidence interval of the predicted AUC scores. We find that the prediction is already better than chance 8-10 months before a user joins the incel forum.
Figure 7: Incel User Case Study: A timeline illustrating the phases of an incel community member exhibiting changes in beliefs towards extremism. Five months prior to joining the community, the user exhibited mild anxiety and frustration with dating. Shortly after joining, they show signs of resentment over rejection and envy, while implying a perceived benefit of joining (to help rid some misery). At 3-4 months, they begin to blame their circumstances on society and use increasingly dehumanizing language about women; note the progression from girl, females, to femoids. At 6-10 months, the user’s rhetoric becomes retributive and violent. The spans are paraphrased extracts from the user’s posts.
Using the same dataset, we additionally examined whether engaging in such communities amplifies extremist traits. We perform local polynomial regression (LOESS) on the extremism scores for each of the users joining the incel community over a period and the users joining a random community over a period of 24 months to derive smooth trajectories of extremism scores. Figure 5 compares the extremism scores of users in incel forums to those in general Reddit forums, analyzing changes in scores before and after joining a new group. For users joining a random, non-extremist forum, we found no significant change in their extremist trait scores. However, for those joining the incel community, we observed a noticeable increase in these traits leading up to and following their entry into the forum; Fig. 7 analyzes the progression of one user’s extremism over this timeline. Despite this, the increase was not as dramatic as one might expect, which aligns with recent research questioning the validity of the “echo chamber” metaphor in explaining polarization and radicalization (e.g., Törnberg and Törnberg 2024). This suggests that joining an extremist group may not automatically intensify extremist tendencies, challenging the assumption that such communities solely function as amplifiers of radicalization.
@@ 9 Limitations and Future Work
Acknowledgements
We would like to express our gratitude to the researchers and scholars whose work has laid the foundation for this study. Their thoughtful and diligent efforts in exploring the complexities of extremism, social psychology, and computational analysis have been invaluable. We would like to thank Isabelle van der Vegt for providing the Stormfront data and making it available to the research team.
This work was supported in part by a grant from the NIH-NIAAA (R01 AA028032) and a DARPA Young Faculty Award grant #W911NF-20-1-0306 awarded to H. Andrew Schwartz at Stony Brook University. The conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of DARPA, NIH, any other government organization, or the U.S. Government. Lucie Flek and Allison Lahnala were supported by the German Federal Ministry of Education and Research (BMBF) as a part of the AI Research Group program under the reference 01-S20060, by the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence and by the Bonn-Aachen International Center for Information Technology (b-it) supporting the visiting research stay at Stony Brook University.
Appendix
Appendix A Term Definitions and Disambiguation
Extremism is the promotion or advancement of an ideology based on violence, hatred, or intolerance that aims to negate fundamental rights, undermine democratic systems, or create permissive environments for such actions (Government of United Kingdom 2024). It represents views and behaviors that are not just unconventional, but rather far removed from mainstream societal norms and attitudes, advocating for extreme measures that threaten democratic systems and the ability of people to live equally under the law. The manifestation of extremist beliefs spans a spectrum, ranging from subtle expressions of intolerance to overt displays of hatred, ultimately culminating in explicit acts of violence against marginalized groups (Schmid 2022; Knight, Woodward, and Lancaster 2017).
Radicalization is the process of change by which an individual or group comes to adopt increasingly extreme views in opposition to a political, social, economic, or religious status quo (Police 2009). This could result in explicit outcomes such as violence or hate-mongering, leading to the degeneration of the individual or the group into extremist ideologies. Radicalization and extremist tendencies usually go hand-in-hand. However, while radicalization focuses more on developing beliefs advocating for a drastic transformation of the social structure, tending towards more and more extreme measures, extremism comprises not just the ideologies but intentional actions and behaviors to propagate those ideologies as well (Trip et al. 2019).
Hate speech is public speech or texts that encourage violence against specific social groups based on race, ethnicity, religion, etc. (Dictionary 2024). While hate speech is commonly studied in NLP contexts (MacAvaney et al. 2019), studying extremism and radicalization online goes beyond just looking for hate speech or speech inciting violence. Although hate speech may stem from deep-rooted extremist beliefs (Bilewicz and Soral 2020), it can also manifest from targeted hostility toward individuals based on personal grievances unrelated to their social identity. (Fortuna, Soler, and Wanner 2020; Piot, Mart´ın-Rodilla, and Parapar 2024) This distinction highlights that hate speech, while potentially overlapping with extremist ideologies and radicalization processes, does not necessarily indicate their presence. Our method is comprehensive – it goes beyond language-based violence incitement to examine fundamental aspects such as underlying beliefs, documented belief transitions, personality traits, dogmatic tendencies, and political, religious, and economic orientations, all grounded in extensive academic research to accurately identify extremist inclinations and radicalization patterns.
Appendix B Banned Subreddit Case Study
Banned r/Right_Wing_Politics r/europeannationalism r/Identitarians r/DebateAltRight r/AltRightChristian r/pol r/NationalSocialism r/AgainstTheChimpire Not-Banned r/AmericanPolitics r/Classical_Liberals r/gunpolitics r/CanadianPolitics r/JordanPeterson r/justicedemocrats r/StillSandersForPres r/CapitalismVSocialism
Appendix C Compute and Hyperparameters
@@
The old-fashioned ways and old-fashioned values still show the best way to live. 0.72 Right-Wing Authoritarianism
@@
God’s laws about abortion, pornography and marriage must be strictly followed before it is too late, violations must be punished. 0.79 Right-Wing Authoritarianism
@@
God’s laws about abortion, pornography and marriage must be strictly followed before it is too late, violations must be punished. 0.79 Right-Wing Authoritarianism
@@
Table 4: Factors of the “Extremist Eleven” along with each factor’s highest scoring extremism scale items.